# HoTPP Benchmark: Are We Good at the Long Horizon Events Forecasting?

### v1

@@ 5 HoTPP Benchmark
The HoTPP benchmark integrates data preprocessing, training, and evaluation in a single toolbox. Unlike previous benchmarks, HoTPP introduces the novel T-mAP metric for long-horizon prediction. HoTPP differs from prior MTPP benchmarks by including simple statistical baselines and next-k models, which simultaneously predict multiple future events. The HoTPP benchmark is designed to focus on simplicity, extensibility, evaluation stability, and reproducibility.
@@
Simplicity and Extensibility. The benchmark code is clearly structured, separating the library, dataset-specific scripts, and configuration files. New methods can be implemented at multiple levels: configuration file, model architecture, loss function, metric, and training module. Core components can be reused by applying Hydra configuration library [Yadan2019Hydra].
Evaluation Stability. Unlike previous works [xue2023easytpp, xue2022hypro], we evaluate methods not only at the end of each sequence but also at intermediate points. This approach collects a larger set of predictions, reducing the final metric’s STD. This is particularly important for datasets with few sequences, such as StackOverflow and Amazon.
@@
Reproducibility. The HoTPP benchmark ensures reproducibility at multiple levels. First, we use the PytorchLightning library [Falcon_PyTorch_Lightning_2019] for training with a specified random seed, and report multi-seed evaluation results. Second, data preprocessing is carefully designed to ensure datasets are constructed reproducibly. Finally, we specify the environment in a Dockerfile.
@@
A detailed benchmark description can be found in Appendix C.
@@
Table 1: Evaluation results. The best result is shown in bold. Mean and STD values of 5 runs with different random seeds are reported.
@@
Method Mean length Next-item Acc., % Next-item MAE OTD Val / Test T-mAP, % Val / Test Transactions MostPopular 7.5 32.78$\pm$ 0.00 0.752$\pm$ 0.000 7.37$\pm$ 0.00 / 7.38$\pm$ 0.00 1.09$\pm$ 0.00 / 1.02$\pm$ 0.00 Last 5 4.7 19.60$\pm$ 0.00 0.924$\pm$ 0.000 7.51$\pm$ 0.00 / 7.55$\pm$ 0.00 1.64$\pm$ 0.00 / 1.69$\pm$ 0.00 MAE-CE 14.0 38.18$\pm$ 0.03 0.570$\pm$ 0.001 7.18$\pm$ 0.01 / 7.20$\pm$ 0.01 4.63$\pm$ 0.33 / 4.50$\pm$ 0.27 MAE-CE-K 9.3 37.68$\pm$ 0.01 0.573$\pm$ 0.000 7.10$\pm$ 0.02 / 7.10$\pm$ 0.02 3.34$\pm$ 0.16 / 3.73$\pm$ 0.16 RMTPP 11.5 38.29$\pm$ 0.02 0.658$\pm$ 0.001 6.78$\pm$ 0.01 / 6.80$\pm$ 0.01 6.72$\pm$ 0.35 / 6.49$\pm$ 0.17 RMTPP-K 9.2 37.74$\pm$ 0.03 0.664$\pm$ 0.001 7.11$\pm$ 0.00 / 7.12$\pm$ 0.00 4.79$\pm$ 0.23 / 4.96$\pm$ 0.13 NHP 4.7 30.49$\pm$ x.xx 0.995$\pm$ x.xxx 7.68$\pm$ x.xx / 7.69$\pm$ x.xx 1.62$\pm$ x.xx / 1.51$\pm$ x.xx HYPRO 11.4 38.27$\pm$ x.xx 0.658$\pm$ x.xxx 6.81$\pm$ x.xx / 6.82$\pm$ x.xx 6.72$\pm$ x.xx / 6.51$\pm$ x.xx MIMIC-IV MostPopular 8.3 1.79$\pm$ 0.00 11.23$\pm$ 0.00 19.89$\pm$ 0.00 / 19.88$\pm$ 0.00 0.73$\pm$ 0.00 / 0.75$\pm$ 0.00 Last 5 4.2 0.88$\pm$ 0.00 4.77$\pm$ 0.00 19.77$\pm$ 0.00 / 19.74$\pm$ 0.00 2.31$\pm$ 0.00 / 2.38$\pm$ 0.00 MAE-CE 10.8 57.38$\pm$ 0.05 2.85$\pm$ 0.01 11.69$\pm$ 0.00 / 11.70$\pm$ 0.01 23.48$\pm$ 0.20 / 23.60$\pm$ 0.20 MAE-CE-K 15.3 56.71$\pm$ 0.22 2.89$\pm$ 0.02 14.04$\pm$ 0.24 / 14.01$\pm$ 0.25 21.26$\pm$ 0.16 / 21.03$\pm$ 0.11 RMTPP 15.9 57.35$\pm$ 0.05 2.85$\pm$ 0.01 12.93$\pm$ 0.03 / 12.93$\pm$ 0.03 20.61$\pm$ 0.15 / 20.85$\pm$ 0.23 RMTPP-K 16.0 56.69$\pm$ 0.03 2.79$\pm$ 0.00 13.91$\pm$ 0.04 / 13.88$\pm$ 0.03 23.01$\pm$ 0.04 / 22.85$\pm$ 0.06 NHP 8.5 24.71$\pm$ 0.95 6.72$\pm$ 0.35 18.73$\pm$ 0.19 / 18.69$\pm$ 0.20 8.55$\pm$ 1.76 / 8.57$\pm$ 1.75 HYPRO 16.0 57.18$\pm$ x.xx 2.83$\pm$ x.xx 12.93$\pm$ x.xx / 12.94$\pm$ x.xx 21.47$\pm$ x.xx / 21.50$\pm$ x.xx Retweet MostPopular 15.5 58.50$\pm$ 0.00 18.82$\pm$ 0.00 174.9$\pm$ 0.0 / 173.5$\pm$ 0.0 26.51$\pm$ 0.00 / 25.87$\pm$ 0.00 Last 10 9.8 50.29$\pm$ 0.00 21.93$\pm$ 0.00 165.5$\pm$ 0.0 / 164.0$\pm$ 0.0 28.54$\pm$ 0.00 / 28.61$\pm$ 0.00 MAE-CE 15.0 59.92$\pm$ 0.04 18.28$\pm$ 0.03 173.6$\pm$ 4.3 / 173.2$\pm$ 4.6 33.10$\pm$ 4.84 / 31.35$\pm$ 4.71 MAE-CE-K 14.8 59.30$\pm$ 0.33 18.25$\pm$ 0.45 169.7$\pm$ 2.2 / 169.3$\pm$ 2.6 36.34$\pm$ 5.06 / 34.09$\pm$ 5.46 RMTPP 12.9 60.06$\pm$ 0.04 18.54$\pm$ 0.09 169.1$\pm$ 2.2 / 167.9$\pm$ 2.3 50.98$\pm$ 0.88 / 48.35$\pm$ 1.01 RMTPP-K 13.6 60.03$\pm$ 0.05 18.42$\pm$ 0.21 166.4$\pm$ 4.7 / 165.8$\pm$ 4.7 49.21$\pm$ 1.06 / 46.31$\pm$ 1.25 NHP 10.7 59.84$\pm$ 0.16 27.37$\pm$ 6.91 194.4$\pm$ 14.6 / 192.1$\pm$ 13.8 39.71$\pm$ 9.02 / 37.52$\pm$ 8.56 HYPRO 13.0 60.12$\pm$ x.xx 18.55$\pm$ x.xx 171.6$\pm$ x.x / 170.6$\pm$ x.x 51.79$\pm$ x.xx / 49.60$\pm$ x.xx Amazon MostPopular 15.7 33.46$\pm$ 0.00 0.304$\pm$ 0.000 7.20$\pm$ 0.00 / 7.18$\pm$ 0.00 9.34$\pm$ 0.00 / 9.02$\pm$ 0.00 Last 5 5.0 24.23$\pm$ 0.00 0.323$\pm$ 0.000 6.70$\pm$ 0.00 / 6.67$\pm$ 0.00 9.73$\pm$ 0.00 / 9.21$\pm$ 0.00 MAE-CE 13.2 35.73$\pm$ 0.12 0.242$\pm$ 0.003 6.62$\pm$ 0.06 / 6.55$\pm$ 0.04 21.98$\pm$ 0.36 / 22.58$\pm$ 0.37 MAE-CE-K 13.3 35.10$\pm$ 0.08 0.246$\pm$ 0.000 6.74$\pm$ 0.01 / 6.70$\pm$ 0.01 21.84$\pm$ 0.05 / 22.35$\pm$ 0.07 RMTPP 15.3 35.74$\pm$ 0.04 0.294$\pm$ 0.001 6.68$\pm$ 0.02 / 6.62$\pm$ 0.03 19.92$\pm$ 0.26 / 20.33$\pm$ 0.33 RMTPP-K 15.8 35.10$\pm$ 0.13 0.300$\pm$ 0.001 6.95$\pm$ 0.01 / 6.89$\pm$ 0.01 17.58$\pm$ 0.12 / 17.83$\pm$ 0.13 NHP 14.4 22.38$\pm$ 6.15 0.378$\pm$ 0.073 7.68$\pm$ 0.12 / 7.67$\pm$ 0.14 21.85$\pm$ 0.86 / 21.34$\pm$ 0.86 HYPRO 14.9 35.87$\pm$ x.xx 0.293$\pm$ x.xxx 6.69$\pm$ x.xx / 6.64$\pm$ x.xx 20.40$\pm$ x.xx / 20.96$\pm$ x.xx StackOverflow MostPopular 11.8 42.90$\pm$ 0.00 0.744$\pm$ 0.000 13.56$\pm$ 0.00 / 13.77$\pm$ 0.00 6.33$\pm$ 0.00 / 5.86$\pm$ 0.00 Last 10 9.3 26.42$\pm$ 0.00 0.950$\pm$ 0.000 14.87$\pm$ 0.00 / 14.92$\pm$ 0.00 8.61$\pm$ 0.00 / 6.64$\pm$ 0.00 MAE-CE 15.7 45.39$\pm$ 0.11 0.641$\pm$ 0.002 13.58$\pm$ 0.08 / 13.65$\pm$ 0.05 8.82$\pm$ 0.70 / 8.45$\pm$ 0.38 MAE-CE-K 15.3 44.84$\pm$ 0.24 0.644$\pm$ 0.003 13.41$\pm$ 0.07 / 13.52$\pm$ 0.06 12.15$\pm$ 0.99 / 11.18$\pm$ 0.79 RMTPP 12.5 45.44$\pm$ 0.13 0.700$\pm$ 0.007 12.97$\pm$ 0.03 / 13.18$\pm$ 0.05 13.44$\pm$ 0.23 / 12.95$\pm$ 0.11 RMTPP-K 12.0 44.84$\pm$ 0.08 0.687$\pm$ 0.001 12.93$\pm$ 0.01 / 13.13$\pm$ 0.02 14.79$\pm$ 0.45 / 14.19$\pm$ 0.26 NHP 10.4 44.44$\pm$ 0.06 0.758$\pm$ 0.021 13.32$\pm$ 0.16 / 13.59$\pm$ 0.16 12.05$\pm$ 0.64 / 11.35$\pm$ 0.42 HYPRO 11.8 45.42$\pm$ x.xx 0.716$\pm$ x.xxx 13.05$\pm$ x.xx / 13.25$\pm$ x.xx 13.51$\pm$ x.xx / 13.61$\pm$ x.xx

### v2

@@ 5 HoTPP Benchmark
The HoTPP benchmark integrates data preprocessing, training, and evaluation in a single toolbox. Unlike previous benchmarks, HoTPP introduces the novel T-mAP metric for long-horizon prediction. HoTPP differs from prior MTPP benchmarks by including simple rule-based baselines and next-k models, simultaneously predicting multiple future events. The HoTPP benchmark is designed with a focus on simplicity, extensibility, evaluation stability, reproducibility, and computational efficiency.
@@
Simplicity and Extensibility. The benchmark code is organized into a clear structure, separating the core library, dataset-specific scripts, and configuration files. New methods can be integrated at various levels, including the configuration file, model architecture, loss function, metric, and training module. Core components are reusable through the Hydra configuration library (Yadan, 2019). For example, HoTPP can apply the HYPRO rescoring method to backbone models such as IFTPP, RMTPP, NHP, and ODE. Another example is our Next-K implementation, which can be applied to IFTTP or RMTPP.
Evaluation Stability. In the MTPP domain, many datasets contain only a few thousand sequences for training and evaluation. For example, the StackOverflow test set includes 401 sequences, Retweet contains 1956, and Amazon has 1851 sequences for testing. Previous long-horizon evaluation pipelines (Xue et al., 2022; 2023) made predictions only at the end of each sequence, resulting in a limited number of predictions and reduced evaluation stability. To address this limitation, we evaluate long-horizon predictions at multiple intermediate points.
@@
Reproducibility. The HoTPP benchmark ensures reproducibility at multiple levels. First, we use the PytorchLightning library (Falcon & The PyTorch Lightning team, 2019) for training with a specified random seed and report multi-seed evaluation results. Second, data preprocessing is carefully designed to ensure datasets are constructed reproducibly. Finally, we specify the environment in a Dockerfile.
Computational Efficiency. Some methods are particularly slow, especially during autoregressive inference on large datasets. Straightforward evaluation can take several hours on a single Nvidia V100 GPU. To accelerate computation, we optimized the training and inference pipelines in two ways. First, we implemented an efficient RNN that reuses computations during parallel autoregressive inference from multiple starting positions. Second, we developed highly optimized versions of the thinning algorithm used for sampling in NHP and continuous-time neural architectures (NHP and ODE), achieving up to a 4x performance improvement compared to the official implementations. These optimizations allowed us to conduct the first large-scale evaluations of algorithms such as NHP, AttNHP, ODE, and HYPRO on datasets like Transactions and MIMIC-IV. Additionally, we provide the first open-source CUDA implementation of the batched linear assignment solver, significantly enhancing the applicability of the proposed T-mAP metric.
@@
A detailed description of the benchmark can be found in Appendix C, with further details on HoTPP’s computational performance in Appendix C.5.
@@
Table 1: Evaluation results. The best result is shown in bold. The mean and standard deviation of each metric computed during five runs with different random seeds are reported.
@@
Method Mean length Next-item Long-horizon Acc (%) mAP (%) MAE OTD Val / Test T-mAP Val / Test (%) Transactions MostPopular 7.5 32.78$\pm$ 0.00 0.86$\pm$ 0.00 0.752$\pm$ 0.000 7.37$\pm$ 0.00 / 7.38$\pm$ 0.00 1.06$\pm$ 0.00 / 0.99$\pm$ 0.00 Last 5 4.7 19.60$\pm$ 0.00 0.87$\pm$ 0.00 0.924$\pm$ 0.000 7.40$\pm$ 0.00 / 7.44$\pm$ 0.00 2.46$\pm$ 0.00 / 2.49$\pm$ 0.00 IFTPP 11.0 34.08$\pm$ 0.04 3.47$\pm$ 0.01 0.693$\pm$ 0.000 6.88$\pm$ 0.01 / 6.90$\pm$ 0.01 5.82$\pm$ 0.27 / 5.88$\pm$ 0.13 IFTPP-K 10.1 33.69$\pm$ 0.03 3.25$\pm$ 0.01 0.698$\pm$ 0.001 7.18$\pm$ 0.00 / 7.19$\pm$ 0.00 4.42$\pm$ 0.15 / 4.43$\pm$ 0.16 RMTPP 7.5 34.15$\pm$ 0.07 3.47$\pm$ 0.02 0.749$\pm$ 0.005 6.86$\pm$ 0.01 / 6.88$\pm$ 0.01 7.08$\pm$ 0.16 / 6.69$\pm$ 0.12 RMTPP-K 7.3 33.63$\pm$ 0.07 3.24$\pm$ 0.02 0.749$\pm$ 0.001 7.10$\pm$ 0.00 / 7.11$\pm$ 0.00 5.82$\pm$ 0.05 / 5.52$\pm$ 0.13 NHP 9.4 35.43$\pm$ 0.07 3.41$\pm$ 0.01 0.696$\pm$ 0.002 6.97$\pm$ 0.01 / 6.98$\pm$ 0.01 5.59$\pm$ 0.07 / 5.61$\pm$ 0.05 AttNHP 7.6 31.12$\pm$ x.xx 1.21$\pm$ x.xx 0.717$\pm$ x.xxx 7.52$\pm$ x.xx / 7.50$\pm$ x.xx 1.71$\pm$ x.xx / 1.48$\pm$ x.xx ODE 9.1 35.60$\pm$ 0.06 3.34$\pm$ 0.06 0.695$\pm$ 0.002 6.96$\pm$ 0.01 / 6.97$\pm$ 0.01 5.53$\pm$ 0.08 / 5.52$\pm$ 0.13 HYPRO 6.9 34.26$\pm$ x.xx 3.46$\pm$ x.xx 0.758$\pm$ x.xxx 7.04$\pm$ x.xx / 7.05$\pm$ x.xx 7.79$\pm$ x.xx / 7.05$\pm$ x.xx MIMIC-IV MostPopular 10.3 4.77$\pm$ 0.00 2.75$\pm$ 0.00 14.52$\pm$ 0.00 19.82$\pm$ 0.00 / 19.82$\pm$ 0.00 0.55$\pm$ 0.00 / 0.54$\pm$ 0.00 Last 5 4.0 1.02$\pm$ 0.00 2.63$\pm$ 0.00 5.34$\pm$ 0.00 19.75$\pm$ 0.00 / 19.73$\pm$ 0.00 2.41$\pm$ 0.00 / 2.49$\pm$ 0.00 IFTPP 12.7 58.59$\pm$ 0.02 47.32$\pm$ 1.39 3.00$\pm$ 0.01 11.51$\pm$ 0.01 / 11.53$\pm$ 0.01 21.93$\pm$ 0.21 / 21.67$\pm$ 0.21 IFTPP-K 15.1 57.91$\pm$ 0.13 44.60$\pm$ 0.14 3.07$\pm$ 0.02 13.17$\pm$ 0.05 / 13.18$\pm$ 0.05 22.46$\pm$ 0.03 / 22.30$\pm$ 0.03 RMTPP 10.0 58.33$\pm$ 0.02 46.24$\pm$ 1.18 3.89$\pm$ 0.03 13.64$\pm$ 0.03 / 13.71$\pm$ 0.03 21.49$\pm$ 0.31 / 21.08$\pm$ 0.29 RMTPP-K 8.4 57.48$\pm$ 0.06 43.47$\pm$ 0.11 3.62$\pm$ 0.02 14.68$\pm$ 0.01 / 14.72$\pm$ 0.02 20.70$\pm$ 0.16 / 20.39$\pm$ 0.14 NHP 6.1 24.97$\pm$ 0.94 11.12$\pm$ 0.49 6.53$\pm$ 0.77 18.59$\pm$ 0.18 / 18.60$\pm$ 0.19 7.26$\pm$ 1.35 / 7.32$\pm$ 1.33 AttNHP 5.1 42.75$\pm$ 0.52 28.55$\pm$ 1.06 3.08$\pm$ 0.04 14.66$\pm$ 0.08 / 14.68$\pm$ 0.08 22.64$\pm$ 0.41 / 22.46$\pm$ 0.40 ODE 12.5 43.21$\pm$ 2.30 25.34$\pm$ 3.05 2.93$\pm$ 0.03 14.71$\pm$ 0.34 / 14.74$\pm$ 0.34 15.41$\pm$ 0.21 / 15.18$\pm$ 0.15 HYPRO 7.4 58.35$\pm$ x.xx 45.45$\pm$ x.xx 3.95$\pm$ x.xx 14.82$\pm$ x.xx / 14.87$\pm$ x.xx 16.94$\pm$ x.xx / 16.77$\pm$ x.xx Retweet MostPopular 28.0 58.50$\pm$ 0.00 39.85$\pm$ 0.00 18.82$\pm$ 0.00 174.9$\pm$ 0.0 / 173.5$\pm$ 0.0 25.15$\pm$ 0.00 / 23.91$\pm$ 0.00 Last 10 9.8 50.29$\pm$ 0.00 35.73$\pm$ 0.00 21.87$\pm$ 0.00 152.3$\pm$ 0.0 / 150.3$\pm$ 0.0 29.12$\pm$ 0.00 / 29.24$\pm$ 0.00 IFTPP 26.1 59.95$\pm$ 0.03 46.53$\pm$ 0.04 18.27$\pm$ 0.03 173.3$\pm$ 4.1 / 172.7$\pm$ 4.4 34.90$\pm$ 4.63 / 31.75$\pm$ 4.44 IFTPP-K 14.8 59.55$\pm$ 0.26 45.09$\pm$ 0.62 18.21$\pm$ 0.42 168.6$\pm$ 2.4 / 167.9$\pm$ 2.8 37.11$\pm$ 4.91 / 34.73$\pm$ 5.11 RMTPP 19.0 60.07$\pm$ 0.10 46.81$\pm$ 0.06 18.45$\pm$ 0.16 167.6$\pm$ 3.4 / 166.7$\pm$ 3.3 47.86$\pm$ 0.91 / 44.74$\pm$ 0.94 RMTPP-K 13.7 59.99$\pm$ 0.05 46.34$\pm$ 0.09 18.33$\pm$ 0.13 164.7$\pm$ 2.8 / 163.9$\pm$ 2.8 49.07$\pm$ 1.16 / 46.16$\pm$ 1.32 NHP 18.1 60.08$\pm$ 0.04 46.83$\pm$ 0.11 18.42$\pm$ 0.11 167.0$\pm$ 1.6 / 165.8$\pm$ 1.6 48.31$\pm$ 0.67 / 45.07$\pm$ 0.34 AttNHP 29.5 60.03$\pm$ 0.03 46.74$\pm$ 0.01 18.39$\pm$ 0.05 173.3$\pm$ 1.0 / 171.6$\pm$ 1.0 28.32$\pm$ 1.24 / 25.85$\pm$ 1.08 ODE 19.0 59.95$\pm$ 0.08 46.65$\pm$ 0.04 18.38$\pm$ 0.04 166.5$\pm$ 0.5 / 165.3$\pm$ 0.5 48.70$\pm$ 0.93 / 44.81$\pm$ 0.69 HYPRO 15.0 59.87$\pm$ x.xx 46.69$\pm$ x.xx 18.75$\pm$ x.xx 171.4$\pm$ x.x / 170.7$\pm$ x.x 49.90$\pm$ x.xx / 46.99$\pm$ x.xx Amazon MostPopular 20.6 33.46$\pm$ 0.00 9.58$\pm$ 0.00 0.304$\pm$ 0.000 7.20$\pm$ 0.00 / 7.18$\pm$ 0.00 8.59$\pm$ 0.00 / 8.31$\pm$ 0.00 Last 5 5.0 24.23$\pm$ 0.00 8.15$\pm$ 0.00 0.321$\pm$ 0.000 6.43$\pm$ 0.00 / 6.41$\pm$ 0.00 9.73$\pm$ 0.00 / 9.21$\pm$ 0.00 IFTPP 13.7 35.73$\pm$ 0.12 17.14$\pm$ 0.10 0.242$\pm$ 0.003 6.58$\pm$ 0.06 / 6.52$\pm$ 0.05 21.94$\pm$ 0.46 / 22.56$\pm$ 0.52 IFTPP-K 13.3 35.11$\pm$ 0.08 16.48$\pm$ 0.03 0.246$\pm$ 0.000 6.72$\pm$ 0.01 / 6.68$\pm$ 0.01 22.06$\pm$ 0.05 / 22.57$\pm$ 0.07 RMTPP 16.7 35.76$\pm$ 0.06 17.21$\pm$ 0.02 0.294$\pm$ 0.001 6.62$\pm$ 0.02 / 6.57$\pm$ 0.03 19.70$\pm$ 0.31 / 20.06$\pm$ 0.33 RMTPP-K 15.8 35.06$\pm$ 0.14 16.37$\pm$ 0.20 0.300$\pm$ 0.005 6.92$\pm$ 0.01 / 6.87$\pm$ 0.02 17.85$\pm$ 0.29 / 18.12$\pm$ 0.30 NHP 9.9 11.06$\pm$ 3.43 11.22$\pm$ 0.19 0.449$\pm$ 0.014 9.04$\pm$ 0.31 / 9.02$\pm$ 0.35 26.24$\pm$ 0.36 / 26.29$\pm$ 0.55 AttNHP 18.5 31.83$\pm$ 0.32 9.70$\pm$ 0.28 0.461$\pm$ 0.003 7.32$\pm$ 0.06 / 7.30$\pm$ 0.06 14.50$\pm$ 0.58 / 14.62$\pm$ 0.80 ODE 9.6 7.54$\pm$ 0.95 10.14$\pm$ 0.23 0.492$\pm$ 0.018 9.48$\pm$ 0.11 / 9.46$\pm$ 0.08 23.54$\pm$ 0.62 / 22.96$\pm$ 0.61 HYPRO 18.0 35.69$\pm$ x.xx 17.21$\pm$ x.xx 0.295$\pm$ x.xxx 6.63$\pm$ x.xx / 6.61$\pm$ x.xx 20.58$\pm$ x.xx / 20.53$\pm$ x.xx StackOverflow MostPopular 14.0 42.90$\pm$ 0.00 5.45$\pm$ 0.00 0.744$\pm$ 0.000 13.56$\pm$ 0.00 / 13.77$\pm$ 0.00 6.10$\pm$ 0.00 / 5.56$\pm$ 0.00 Last 10 9.3 26.42$\pm$ 0.00 5.20$\pm$ 0.00 0.934$\pm$ 0.000 14.52$\pm$ 0.00 / 14.55$\pm$ 0.00 8.67$\pm$ 0.00 / 6.72$\pm$ 0.00 IFTPP 24.1 45.41$\pm$ 0.11 13.00$\pm$ 0.79 0.641$\pm$ 0.002 13.57$\pm$ 0.07 / 13.64$\pm$ 0.05 8.78$\pm$ 0.87 / 8.31$\pm$ 0.50 IFTPP-K 15.3 44.85$\pm$ 0.24 11.16$\pm$ 1.07 0.644$\pm$ 0.003 13.41$\pm$ 0.07 / 13.51$\pm$ 0.06 12.42$\pm$ 0.97 / 11.42$\pm$ 0.78 RMTPP 14.5 45.43$\pm$ 0.13 13.33$\pm$ 0.25 0.701$\pm$ 0.007 12.95$\pm$ 0.02 / 13.17$\pm$ 0.05 13.26$\pm$ 0.29 / 12.72$\pm$ 0.16 RMTPP-K 12.0 44.89$\pm$ 0.09 11.72$\pm$ 0.10 0.689$\pm$ 0.001 12.92$\pm$ 0.01 / 13.13$\pm$ 0.01 14.91$\pm$ 0.19 / 14.30$\pm$ 0.09 NHP 13.0 44.53$\pm$ 0.05 10.86$\pm$ 0.19 0.715$\pm$ 0.004 13.02$\pm$ 0.02 / 13.24$\pm$ 0.02 12.67$\pm$ 0.55 / 11.96$\pm$ 0.40 AttNHP 15.5 45.17$\pm$ 0.10 12.67$\pm$ 0.13 0.705$\pm$ 0.001 13.08$\pm$ 0.03 / 13.30$\pm$ 0.02 11.95$\pm$ 0.19 / 11.13$\pm$ 0.32 ODE 13.9 44.38$\pm$ 0.09 10.12$\pm$ 0.19 0.711$\pm$ 0.004 13.04$\pm$ 0.02 / 13.27$\pm$ 0.03 11.37$\pm$ 0.32 / 10.52$\pm$ 0.23 HYPRO 11.8 45.18$\pm$ x.xx 12.88$\pm$ x.xx 0.715$\pm$ x.xxx 13.04$\pm$ x.xx / 13.26$\pm$ x.xx 15.57$\pm$ x.xx / 14.69$\pm$ x.xx